# aiXcoder-7B: A Lightweight and Effective Large Language Model for Code Completion

### v2

@@ Abstract
Large Language Models (LLMs) have been widely used in code completion, and researchers are focusing on scaling up LLMs to improve their accuracy. However, larger LLMs will increase the response time of code completion and decrease the developers’ productivity. In this paper, we propose a lightweight and effective LLM for code completion named aiXcoder-7B. Compared to existing LLMs, aiXcoder-7B achieves higher code completion accuracy while having smaller scales (i.e., 7 billion parameters). We attribute the superiority of aiXcoder-7B to three key factors: ❶ Multi-objective training. We employ three training objectives, one of which is our proposed Structured Fill-In-the-Middle (SFIM). SFIM considers the syntax structures in code and effectively improves the performance of LLMs for code. ❷ Diverse data sampling strategies. They consider inter-file relationships and enhance the capability of LLMs in understanding cross-file contexts. ❸ Extensive high-quality data. We establish a rigorous data collection pipeline and consume a total of 1.2 trillion unique tokens for training aiXcoder-7B. This vast volume of data enables aiXcoder-7B to learn a broad distribution of code. We evaluate aiXcoder-7B in five popular code completion benchmarks and a new benchmark collected by this paper. The results show that aiXcoder-7B outperforms the latest six LLMs with similar sizes and even surpasses four larger LLMs (e.g., StarCoder2-15B and CodeLlama-34B), positioning aiXcoder-7B as a lightweight and effective LLM for academia and industry. Finally, we summarize three valuable insights for helping practitioners train the next generations of LLMs for code. aiXcoder-7B has been open-souced and gained significant attention [1]. As of the submission date, aiXcoder-7B has received 2,193 GitHub Stars.
@@ I Introduction
We assess the effectiveness of aiXcoder-7B in three code completion tasks, including Fill-In-the-Middle (FIM), cross-file code completion, and Natural language to Code (NL2Code). We experiment with six code completion benchmarks, five of which are popular public datasets and one is FIM-Eval collected by this paper. FIM-Eval is a benchmark for FIM, consisting of 16,136 samples and covering four languages (i.e., Java, Python, C++, JavaScript). FIM-Eval additionally labels the types of code to be completed, including 13 types, e.g., function signatures and comments. Then, we compare aiXcoder-7B to 10 recently released LLMs (from 7B to 34B) on these benchmarks and yield the following insights: ❶ aiXcoder-7B substantially outperforms LLMs with similar sizes and even surpasses larger LLMs in six benchmarks. For example, in a popular benchmark - HumanEval, aiXcoder-7B achieves a Pass@1 score of 54.9%, outperforming CodeLlama-34B (i.e., 48.2%) and StarCoder2-15B (i.e., 46.3%). The improvements show that aiXcoder-7B achieves higher code completion accuracy while having smaller scales. ❷ Based on our FIM-Eval, we analyze the performance of aiXcoder-7B in completing different types of code. aiXcoder-7B outperforms LLMs with similar sizes on most types (max: 13, min: 8). The results show the strong generalization ability of aiXcoder-7B in code completion. ❸ We show that existing LLMs are prone to generate longer code in FIM, while the code generated by aiXcoder-7B is closer in length to human-written reference code. The result shows that the code generated by aiXcoder-7B is more concise and closer to the human coding style.
@@
  • We evaluate the effectiveness of aiXcoder-7B in six code completion benchmarks. aiXcoder-7B substantially outperforms 6 LLMs with similar sizes and even surpasses 4 larger LLMs (15B and 34B).
@@ II-C Data Deduplication
  • Exact deduplication. We extract file contents, find the files with exactly the same content, and keep only one copy.
@@ II-E Sensitive Information Removal
Match-based filter. We manually build a list of sensitive words, which covers a broad range of sensitive topics (e.g., politics). Then, we scan all pre-training data and delete the data containing sensitive words.
@@ III-C Training Objectives
Next-Token Prediction (NTP). It is similar to code completion, training models to predict the subsequent token based on the provided context. Given a code file or natural language article $\mathbf{x}=\{x_{0},x_{1},\dots,x_{l}\}$, NTP trains models predict the next token $x_{i}$ based on previous tokens $\{x_{<i}\}$. The objective is to minimize the following loss function:
@@ IV-A Research Questions
RQ2: How does aiXcoder-7B perform on Fill-In-the-Middle task compared to existing LLMs? FIM simulate scenarios where developers modify existing code by predicting the missing middle portion using bidirectional contexts.
@@
In the RQs, we apply aiXcoder-7B on 6 benchmarks totally. These benchmarks cover 6 programming languages. To show the superiority of aiXcoder-7B, we also select 10 popular LLMs as baselines for comparison. Then, we report the execution-based and text-based metrics (Section IV-D) of completed programs.
@@ IV-B Compared Models
Note that our aiXcoder-7B was open-sourced in March 2024. Thus, we select LLMs released before March 2024 for comparison. Specifically, we select 10 popular LLMs for comparison, and they can be divided into two groups:
@@
  • DeepSeekCoder-7B [3], by DeepSeek AI, is a 7B parameter model trained on a blend of code and natural language data, designed for programming tasks.
@@ IV-C Benchmarks
  • FIM-Eval is a large-scale FIM benchmark collected by this paper. We construct FIM-Eval from some real-world repositories, which are excluded from the training data of aiXcoder-7B. We extract 13 types of code snippets from these repositories and randomly mine spans from these code snippets. These 13 types of code snippets encompass common code completion scenarios, including method signatures, method bodies, single-line statements, methods with comments, empty code blocks, specific positions within a method body (top, middle, and bottom), specific control statements (i.e., if statements, for loops, while loops, try statements, and switch-case statements). Finally, we collect 16,140 samples covering four programming languages: C++ (4,080 samples), Java (4,080 samples), Python (3,900 samples), and JavaScript (4,080 samples). FIM-Eval provides a reliable, practical, and diverse evaluation platform for FIM. FIM-Eval has been open-sourced in our repository [1].
@@
  • CrossCodeEval [25] covers four popular programming languages: 2,665 Python samples, 2,139 Java samples, 3,356 TypeScript samples, and 1,768 C# samples. Each sample is provided in three formats: no cross-file context, retrieved cross-file context, and retrieval with reference. The LLMs completed code snippet is compared using text-based metrics.
@@ V-C RQ3: Performance on Cross-File Code Completion
Another important capability of LLMs is the ability to understand code context across files, as developers often need to consider information from other files within the current project when writing code. In Table IV, we fix the context length for all LLMs at 16K and format the input using the PSM pattern in FIM. All LLMs employ the greedy search to generate code.
We design three experimental settings: ❶ Base Setting. As a baseline, LLMs complete based solely on the current file without cross-fire context. ❷ Retrieval BM25. Based on the current file context in the base settings, It additionally uses BM25 to match repository code fragments. The top 5 matches, capped at 512 tokens, are added to the prompt, along with formatted class definitions from other files. ❸ Retrieval w/Ref. In this setting, we make use of not only the in-file context (as in Retrieval BM25 setting) but also the reference to retrieve the cross-file context. We prepend the retrieved context to the in-file context to construct the prompt for this setting.
@@ VI Discussion
VI-A Comparison in the Length of Code
@@
To gain insights into this aspect, we evaluate LLMs’ performance using FIM-Eval (Section IV-C), which includes a variety of scenarios. Additionally, we present the code length ratio, which is calculated as the ratio of the number of tokens in the prediction to the number of tokens in the ground truth code. Based on the experimental results in Table VI below, we observe that existing LLMs tend to over-generate, producing code that is substantially longer than necessary. Too long code will increase the burden on users and reduce maintainability. For example, CodeLlama-7B produces ratios of 2.14 for Java and 3.02 for C++, while StarCoder2-7B reaches even higher ratios, such as 3.62 for C++. In contrast, aiXcoder-7B consistently generates code predictions that are similar in length to human-written code, achieving ratios of 0.97 for Java and 0.87 for Python. We attribute this performance to aiXcoder-7B’s Structured Fill-In-the-Middle (SFIM) training objective, which helps align model outputs with human-written code, resulting in more efficient coding practices.
TABLE VI: The length comparison between the generated code and reference code
@@
VI-B Insights of Training LLMs for Code
@@
Scaling up training data can continuously improve the performance of LLMs. Although the scaling law [29] provides a relationship between model size and the amount of training data, we discover that further scaling up training data is necessary. Even if the training loss of LLMs is already small, continually training with new data still can improve the performance of models. Similar phenomena have also been found in other works [30].

### v3

# aiXcoder-7B: A Lightweight and Effective Large Language Model for Code Processing

@@ Abstract
Large Language Models (LLMs) have been widely used in code completion, and researchers are focusing on scaling up LLMs to improve their accuracy. However, larger LLMs have lower inference efficiency, affecting developers’ experience and productivity. In this paper, we propose a lightweight and effective LLM for code completion named aiXcoder-7B. Compared to existing LLMs, aiXcoder-7B achieves higher code completion accuracy while having smaller scales (i.e., 7 billion parameters). We attribute the superiority of aiXcoder-7B to three key factors: ❶ Multi-objective training. We employ three training objectives, one of which is our proposed Structured Fill-In-the-Middle (SFIM). SFIM considers the syntax structures in code and effectively improves the performance of LLMs for code. ❷ Diverse data sampling strategies. They consider inter-file relationships and enhance the capability of LLMs in understanding cross-file contexts. ❸ Extensive high-quality data. We establish a rigorous data collection pipeline and consume a total of 1.2 trillion unique tokens for training aiXcoder-7B. This vast volume of data enables aiXcoder-7B to learn a broad code distribution. We evaluate aiXcoder-7B in five popular code completion benchmarks and a new benchmark collected by this paper. The results show that aiXcoder-7B outperforms the latest seven LLMs with similar sizes and even surpasses four larger LLMs (e.g., StarCoder2-15B and CodeLlama-34B), positioning aiXcoder-7B as a lightweight and effective LLM for academia and industry. Finally, we summarize three valuable insights for helping practitioners train the next generations of LLMs for code. aiXcoder-7B has been open-sourced and gained significant attention [1]. Until January 2025, aiXcoder-7B has received 2,226 GitHub Stars.
@@ I Introduction
We assess the effectiveness of aiXcoder-7B in three code completion tasks, including Fill-In-the-Middle (FIM), cross-file code completion, and Natural language to Code (NL2Code). We experiment with six code completion benchmarks, five of which are popular public datasets and one is FIM-Eval collected by this paper. FIM-Eval is a benchmark for FIM, consisting of 16,136 samples and covering four languages (i.e., Java, Python, C++, JavaScript). FIM-Eval additionally labels the types of code to be completed, including 13 types, e.g., function signatures and comments. Then, we compare aiXcoder-7B to 11 recently released LLMs (from 7B to 34B) on these benchmarks and yield the following insights: ❶ aiXcoder-7B substantially outperforms LLMs with similar sizes and even surpasses larger LLMs in six benchmarks. For example, in a popular benchmark - HumanEval, aiXcoder-7B achieves a Pass@1 score of 54.9%, outperforming CodeLlama-34B (i.e., 48.2%) and StarCoder2-15B (i.e., 46.3%). The improvements show that aiXcoder-7B achieves higher code completion accuracy while having smaller scales. ❷ Based on our FIM-Eval, we analyze the performance of aiXcoder-7B in completing different types of code. aiXcoder-7B outperforms LLMs with similar sizes on most types (max: 13, min: 8). The results show the strong generalization ability of aiXcoder-7B in code completion. ❸ We show that existing LLMs are prone to generate longer code in FIM, while the code generated by aiXcoder-7B is closer in length to human-written reference code. The result shows that the code generated by aiXcoder-7B is more concise and closer to the human coding style.
@@
  • We evaluate the effectiveness of aiXcoder-7B in six code completion benchmarks. aiXcoder-7B substantially outperforms 7 LLMs with similar sizes and even surpasses 4 larger LLMs (15B and 34B).
Paper Organization. Section II presents our data collection pipeline. Section III describes the training process of aiXcoder-7B. Section IV and Section V show the study design and results, respectively. Section VI discusses the performance of aiXcoder-7B in other aspects and potential threats. Section VII summarizes the recently proposed LLMs. Finally, Section VIII concludes this paper and points out future directions.
@@ II-C Data Deduplication
  • Exact deduplication. We extract file contents, find the files with the same content, and keep only one copy.
@@ II-E Sensitive Information Removal
Match-based filter. We manually build a list of sensitive words, which covers a broad range of sensitive topics (e.g., politics). Then, we scan all pre-training data and delete the data containing sensitive words. For example, if the frequency of sensitive words in a repository of a user exceeds a certain threshold, all repositories of the user will be deleted.
@@ III-C Training Objectives
Next-Token Prediction (NTP). It is similar to code completion, training models to predict the subsequent token based on the provided context. Given a code file or natural language article $\mathbf{x}=\{x_{0},x_{1},\dots,x_{l}\}$, NTP trains models to predict the next token $x_{i}$ based on previous tokens $\{x_{<i}\}$. The objective is to minimize the following loss function:
@@ IV-A Research Questions
RQ2: How does aiXcoder-7B perform on Fill-In-the-Middle task compared to existing LLMs? FIM simulates scenarios where developers modify existing code by predicting the missing middle portion using bidirectional contexts.
@@
In the RQs, we apply aiXcoder-7B on 6 benchmarks in total. These benchmarks cover 6 programming languages. To show the superiority of aiXcoder-7B, we also select 10 popular LLMs as baselines for comparison. Then, we report the execution-based and text-based metrics (Section IV-D) of completed programs.
@@ IV-B Compared Models
We select 10 popular LLMs for comparison, and they can be divided into two groups:
@@
  • DeepSeekCoder-7B [3], by DeepSeek AI, is a 7B parameter model trained on a blend of code and natural language data designed for programming tasks.
  • Qwen2.5-Coder-7B [21] is a large language model with 7 billion parameters. It builds on the strong Qwen2.5 [22] and continues training on a larger scale of code data, including source code, text-code grounding data, and synthetic data, totaling 5.5 trillion tokens. This leads to significant improvements in code-related tasks.
@@ IV-C Benchmarks
  • FIM-Eval is a large-scale FIM benchmark collected by this paper. We construct FIM-Eval from some real-world repositories, which are excluded from the training data of aiXcoder-7B. We extract 13 types of code snippets from these repositories and randomly mine spans from these code snippets. These 13 types of code snippets encompass common code completion scenarios, including method signatures, method bodies, single-line statements, methods with comments, empty code blocks, specific positions within a method body (top, middle, and bottom), specific control statements (i.e., if statements, for loops, while loops, try statements, and switch-case statements). Finally, we collected 16,140 samples covering four programming languages: C++ (4,080 samples), Java (4,080 samples), Python (3,900 samples), and JavaScript (4,080 samples). FIM-Eval provides a reliable, practical, and diverse evaluation platform for FIM. FIM-Eval has been open-sourced in our repository [1].
@@
  • CrossCodeEval [29] covers four popular programming languages: 2,665 Python samples, 2,139 Java samples, 3,356 TypeScript samples, and 1,768 C# samples. Each sample is provided in three formats: no cross-file context, retrieved cross-file context, and retrieval with reference. The completed code snippet is compared using text-based metrics.
@@ V-C RQ3: Performance on Cross-File Code Completion
Another important capability of LLMs is the ability to understand code context across files, as developers often need to consider information from other files within the current project when writing code [33, 34]. In Table IV, we fix the context length for all LLMs at 16K and format the input using the PSM pattern in FIM. All LLMs employ the greedy search to generate code.
We design three experimental settings: ❶ Base Setting. As a baseline, LLMs predict the code based solely on the current file without cross-fire context. ❷ Retrieval BM25. Based on the current file context in the base settings, it additionally uses BM25 to match repository code fragments. The top 5 matches, capped at 512 tokens, are added to the prompt. ❸ Retrieval w/Ref. In this setting, we use not only the in-file context (as in Retrieval BM25 setting) but also the reference to retrieve the cross-file context. We prepend the retrieved context to the in-file context to construct the prompt for this setting.
@@ VI Discussion
VI-A Why choose 7B?
TABLE VI: The inference efficiency (i.e., time taken to generate a completion suggestion) of aiXcoder-7B (bfloat16) and Qwen2.5-Coder-32B (int4).
Input Length Output Length Qwen2.5-Coder-32B aiXcoder-7B
1024 128 23.70 ms 14.81 ms
4096 128 32.71 ms 17.10 ms
8192 128 64.70 ms 21.79 ms
16384 128 91.65 ms 29.26 ms
The inference efficiency of code completion models is crucial for enhancing the developer’s experience. While scaling up model size can improve accuracy, it often comes at the cost of inference efficiency. Table VI illustrates the inference efficiency of our aiXcoder-7B (bfloat16) and Qwen2.5-Coder-32B (int4) under identical testing settings on a single A100 GPU. Despite the quantization of Qwen2.5-Coder-32B to int4, which slightly reduces its inference quality, its inference speed and throughput are still significantly lower compared to aiXcoder-7B in bfloat16. For example, Qwen2.5-Coder-32B requires 91.65 ms to generate a completion suggestion, whereas aiXcoder-7B achieves a much faster response time of 29.26 ms—approximately 3 times faster. These results demonstrate that aiXcoder-7B can generate completion suggestions both accurately and efficiently, making it superior for practical applications.
VI-B The Comparison in the Length of Code
@@
To gain insights into this aspect, we evaluate LLMs’ performance using FIM-Eval (Section IV-C), which includes a variety of scenarios. Additionally, we present the code length ratio, which is calculated as the ratio of the number of tokens in the prediction to the number of tokens in the ground truth code. Based on the experimental results in Table VII below, we observe that existing LLMs tend to over-generate, producing code that is substantially longer than necessary. Too long code will increase the burden on users and reduce maintainability. For example, CodeLlama-7B produces ratios of 2.14 for Java and 3.02 for C++, while StarCoder2-7B reaches even higher ratios, such as 3.62 for C++. In contrast, aiXcoder-7B consistently generates code predictions that are similar in length to human-written code, achieving ratios of 0.97 for Java and 0.87 for Python. We attribute this performance to aiXcoder-7B’s Structured Fill-In-the-Middle (SFIM) training objective, which helps align model outputs with human-written code, resulting in more efficient coding practices.
TABLE VII: The length comparison between the generated code and reference code
@@
VI-C Insights of Training LLMs for Code
@@
Scaling up training data can continuously improve the performance of LLMs. Although the scaling law [35] provides a relationship between model size and the amount of training data, we discover that further scaling up training data is necessary. Even if the training loss of LLMs is already small, continually training with new data can still improve the performance of models. Similar phenomena have also been found in other works [36].
@@ VIII Conclusion and Future Work
Acknowledgment
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the Major Program (JD) of Hubei Province (No.2023BAA024), and the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62072007.